### v1

# Second Law of Thermodynamics without Einstein Relation

Abstract
Materials that are constantly driven out of thermodynamic equilibrium, such as active and living systems, typically violate the Einstein relation. This may arise from active contributions to particle fluctuations which are unrelated to the dissipative resistance of the surrounding medium. We show that in these cases the widely used relation between informatic entropy production and heat dissipation does not hold. Consequently, fluctuation relations for the mechanical work, such as the Jarzynski and Crooks theorems, are invalid. We relate the breaking of the correspondence between entropy production and heat dissipation to the breaking of the fluctuation-dissipation theorem. We propose a temperature-like variable which restores this correspondence and gives rise to a generalized second law of thermodynamics, whereby the dissipated heat is necessarily nonnegative and vanishes at equilibrium. The Clausius inequality, Carnot maximum efficiency theorem, and relation between the maximum extractable work and the change of free energy are recovered as well.
Classical thermodynamics is the central theory for the possible forms of energy transfer in materials. It was historically formulated for the purpose of estimating the work that can be extracted from heat engines, utilizing the principle that heat spontaneously flows from “hotter” to “colder” objects. This is one version of the second law of thermodynamics, which determines the direction of spontaneous processes. Comparing the efficiencies of different heat engines provides a definition of a consistent “hotness” scale which is a thermodynamic state function, i.e., the temperature $T$ [1, 2]. In light of its vast applicability and usefulness for characterizing the behaviors of equilibrium materials, there is an ongoing effort to extend the theory of thermodynamics to systems that are constantly driven out of equilibrium, for example, self-propelled particles [3, 4, 5], chemical reactions [6, 7], and living systems [8]. Among the complications associated with nonequilibrium systems is the inadequacy of temperature and other variables as thermodynamic state functions [2, 9, 10, 11, 12, 13].
One of the consequences of thermodynamic equilibrium, where the system’s configurations are Boltzmann-distributed, is the Einstein relation (ER) [14, 15],
$\mathbf{D}=k_{\mathrm{B}}T\boldsymbol{\mu}\,.$ (1)
The diffusivity $\mathbf{D}$ and mobility $\boldsymbol{\mu}$ characterize, respectively, the strength of thermal fluctuations and the dissipative attenuation of motion. In general they may be space-dependent matrices, reflecting translational and rotational symmetry breaking. At equilibrium the effects of fluctuations and dissipation exactly balance [15], leading to Eq. (1). However, in cases where the forces are not conservative [16, 9], or if the fluctuations arise from coarse-graining of fast athermal or active forces [17, 18, 11, 19, 20, 21, 22], there should not be a generic relation between fluctuations (diffusivity) and dissipation (mobility). Thus, more often than not, in far-from-equilibrium cases such as active biopolymer networks [23, 24, 25], bacterial swarms [26, 27], and chemical reactions out of equilibrium [28, 7], the ER does not hold. In cases where $\mathbf{D}$ and $\boldsymbol{\mu}$ are proportional to one another, one can use Eq. (1) to define an effective temperature [2, 11, 17, 29]. However, in the general case these two position-dependent matrices need not be related by a constant uniform scalar, and Eq. (1) is violated.
One of the theories addressing nonequilibrium systems is stochastic thermodynamics [30]. By considering Langevin equations or discrete Markov chains it has succeeded in establishing trajectory-wise energy conservation (leading to the first law of thermodynamics) [31, 32, 33] and proving a wealth of fluctuation relations (FRs) [34], such as the Jarzynski equality [35, 36] and Crooks relation [37]. The latter also yields the second law of thermodynamics for the average entropy changes and heat increments (with some individual trajectories violating it [38]). Yet, implicitly, all the aforementioned FRs and extended laws of thermodynamics have relied on the existence of a temperature, defined through the ER, to connect the microscopic statistics to thermodynamic energy, mechanical and chemical work, and calorimentric heat.
Several definitions of a generalized nonequilibrium temperature were proposed based on generalizations of the fluctuation-dissipation theorem (FDT) [39, 11, 40, 18]. Most of these definitions still assume a thermal bath that obeys the ER. In other cases the ER is required to keep the extended FRs connected to physical measurables. For example, the Hatano-Sasa relation [41], an extension of the Jarzynski equality, requires Markovian dynamics only. It therefore relaxes Jarzynski’s assumption of a steady-state Boltzmann distribution for a given value of the external protocol [36], thus relaxing also the ER requirement. Yet, the effective Hamiltonian and work used by this relation coincide with their actual, measurable counterparts only under those extra assumptions. Otherwise, they demand knowledge of the complete microstates’ steady-state distribution, which is usually inaccessible or requires simplistic models [42]. The generalized temperature proposed here, arising directly from thermodynamics, does not rely on those assumptions.
We have two aims: (i) to investigate the fate of existing nonequilibrium FRs and extensions to the second law of thermodynamics once the assumption of an ER-based temperature is removed; and (ii) to formulate a generalized, ER-independent temperature and derive the resulting alternative thermodynamic identities.
Overdamped stochastic energetics.—We consider an ensemble of $N$ interacting degrees of freedom (particles), subjected to a deterministic external protocol $\Lambda_{t}$, which is a prechosen time-varying system parameter. The time-dependent stochastic microstate is an $N$-component vector, $\mathbf{X}_{t}$. We couple the particles to an arbitrary medium, which is assumed to provide them with effective mobility $\boldsymbol{\mu}(\mathbf{x},\lambda)$ and diffusivity $\mathbf{D}(\mathbf{x},\lambda)$ tensors, such that the particles follow the overdamped Langevin equation [Eq. (2) below]; no other property of the surrounding medium is specified. The aforementioned $\mathbf{X}_{t}$ and $\Lambda_{t}$ denote a particular time-dependent realizations of $\mathbf{x}$ and $\lambda$. The particles are subjected to an additional external force $\mathbf{F}_{\mathrm{ext}}(\mathbf{x},\lambda)$, which may be nonconservative and is not included in the internal energy. We denote the total ($N$-component) force acting on the particles by $\mathbf{F}=\mathbf{F}_{\mathrm{ext}}-\boldsymbol{\nabla}H$, where $H(\mathbf{x},\lambda)$ is the system’s Hamiltonian. The system is driven out of equilibrium by applying the nonconservative force $\mathbf{F}_{\mathrm{ext}}$ and/or violating the ER [Eq. (1)]. Thus, the system’s steady-state distribution is not Boltzmann, nor can a temperature be conventionally identified.
The system follows the overdamped Langevin equation,
$\dot{\mathbf{X}}_{t}$ $=$ $\boldsymbol{\mu}(\mathbf{X}_{t},\Lambda_{t})\cdot\mathbf{F}(\mathbf{X}_{t},\Lambda_{t})+{\boldsymbol{\nabla}}\cdot\mathbf{D}(\mathbf{X}_{t},\Lambda_{t})$ (2)
$+\mathbf{b}(\mathbf{X}_{t},\Lambda_{t})\cdot\boldsymbol{\eta}_{t}\,,$
where $\mathbf{b}$ is the noise magnitute satisfying $\mathbf{D}=\mathbf{b}\cdot\mathbf{b}^{\mathrm{T}}/2$, $\boldsymbol{\eta}_{t}$ is the normalized thermal noise [$\langle\boldsymbol{\eta}_{t}\rangle=\mathbf{0}$, $\langle\boldsymbol{\eta}_{t}\boldsymbol{\eta}_{t^{\prime}}\rangle=\mathbf{I}\delta(t-t^{\prime})$ where $\mathbf{I}$ is the identity matrix], and the equation is written in the Itô convention [43]. The so-called “spurious drift” term ${\boldsymbol{\nabla}}\cdot\mathbf{D}$ ensures that the Boltzmann distribution is obtained at equilibrium, i.e., in systems with a well-defined temperature, Eq. (1), under the effect of conservative forces [44, 45]. The corresponding Fokker-Planck equation [46] for the microstates distribution $p(\mathbf{x},t)$ is
$\frac{\partial p(\mathbf{x},t)}{\partial t}$ $=$ $-{\boldsymbol{\nabla}}\cdot\mathbf{J}(\mathbf{x},t)\,,$ (3)
$\mathbf{J}$ $=$ $\boldsymbol{\mu}\cdot\mathbf{F}\,p-\mathbf{D}\cdot{\boldsymbol{\nabla}}p\,,$ (4)
where $\mathbf{J}$ is the probability flux.
We now follow Sekimoto’s approach [31, 32] to identify the internal energy, work, and heat for trajectories in generic Langevin systems. The time-derivative of the Hamiltonian is [43]
$\dot{H}_{t}={\boldsymbol{\nabla}}H\circ\dot{\mathbf{X}}_{t}+\frac{\partial H}{\partial\lambda}\dot{\Lambda}_{t}\,,$ (5)
where $\circ$ is the Stratonovich product [47]. The term $(\partial H/\partial\lambda){\text{d}}\Lambda_{t}$ is the external work necessary to change the system’s internal energy over time ${\text{d}}t$. The force $\mathbf{F}_{\mathrm{ext}}(\mathbf{X}_{t},\Lambda_{t})$ acting along a displacement ${\text{d}}\mathbf{X}_{t}$ during ${\text{d}}t$ performs an additional work, $\mathbf{F}_{\mathrm{ext}}\circ{\text{d}}\mathbf{X}_{t}$. Thus the total work rate is
$\dot{W}_{t}=\frac{\partial H}{\partial\lambda}\dot{\Lambda}_{t}+\mathbf{F}_{\mathrm{ext}}\circ\dot{\mathbf{X}}_{t}\,.$ (6)
According to the first law of thermodynamics, the heat rate at time $t$ is the remaining energy flow into the system,
$\dot{Q}_{t}=\dot{H}_{t}-\dot{W}_{t}=-\mathbf{F}\circ\dot{\mathbf{X}}_{t}\,.$ (7)
Converting the Stratonovich differential into Itô [43] and inserting Eq. (2), we find an implicit Langevin equation for the heat in terms of the solutions $\mathbf{X}_{t}$ of Eq. (2),
$\dot{Q}_{t}=-\mathbf{F}\cdot\boldsymbol{\mu}\cdot\mathbf{F}-{\boldsymbol{\nabla}}\cdot(\mathbf{D}\cdot\mathbf{F})-\mathbf{F}\cdot\mathbf{b}\cdot\boldsymbol{\eta}_{t}\,.$ (8)
The above is a natural adaptation of the internal energy, work, and heat to the microscale [31, 32]. A similar extension of the system’s entropy is less trivial. It is common to consider the instantaneous nonequilibrium stochastic entropy [30, 34], $S(\mathbf{x},t)=-\ln p(\mathbf{x},t)$, whose ensemble average is the usual Shannon entropy [48, 49, 50], $\langle S\rangle=-\int{\text{d}}\mathbf{x}\,p(\mathbf{x},t)\ln p(\mathbf{x},t)$. The relation between classical-thermodynamic entropy and stochastic entropy is established for equilibrium states by finding the distribution $p$ that maximizes the Shannon entropy $\langle S\rangle$ [1] (for a given fixed protocol value $\lambda$). Combining the stochastic chain rule [43] and Eqs. (2)–(4), the time-derivative of $S_{t}=S(\mathbf{X}_{t},t)$ is
$\dot{S}_{t}={\boldsymbol{\nabla}}\cdot\left(\frac{\mathbf{J}}{p}\right)-\frac{1}{p}{\boldsymbol{\nabla}}\cdot(\mathbf{D}\cdot{\boldsymbol{\nabla}}p)-[({\boldsymbol{\nabla}}\ln p)\cdot\mathbf{b}]\cdot\boldsymbol{\eta}_{t}\,,$ (9)
where the last two terms are often omitted [51] as they vanish on average.
If a well-defined temperature $T$ exists, it is instructive to consider the thermodynamic heat dissipation $R_{t}$. It quantifies how much extra calorimetric heat $\dot{Q}_{t}$ is wasted in a given irreversible process. According to the second law of thermodynamics, the average thermodynamic dissipation $\langle\dot{R}_{t}\rangle$ is nonnegative [38],
$\dot{R}_{t}\equiv k_{\mathrm{B}}T\dot{S}_{t}-\dot{Q}_{t}\,,\qquad\langle\dot{R}_{t}\rangle\geq 0\,,$ (10)
where $k_{\mathrm{B}}T\dot{S}_{t}$ is the reversible contribution to the heat. Our goal is to arrive at a similar relation in the absence of the conventional temperature $T$.
Informatic entropy production (IEP).—The IEP, $\Sigma_{t}$, quantifies the breaking of time-reversal symmetry in the microstates’ dynamics [30]. We consider a discretized trajectory drawn from Eq. (2), $(\mathbf{X}_{0},\dots,\mathbf{X}_{K\Delta t})=(\mathbf{x}_{0},\dots,\mathbf{x}_{K})$, with an arbitrarily-small timestep $\Delta t$. The probability distribution of obtaining this trajectory under the effect of the protocol sequence $(\lambda_{0},\lambda_{1},\ldots,\lambda_{K})$, $\overrightarrow{\mathcal{P}}\equiv\Pr(\mathbf{x}_{0},\mathbf{x}_{1},\ldots,\mathbf{x}_{K})$, is compared with its reversed counterpart — the probability distribution of obtaining the reverted trajectory, $\overleftarrow{\mathcal{P}}\equiv\Pr(\mathbf{x}_{K},\mathbf{x}_{K-1},\ldots,\mathbf{x}_{0})$, under the reverted protocol $(\lambda_{K},\lambda_{K-1},\ldots,\lambda_{0})$. Explicitly [34, 30],
$\Sigma_{t}=\ln\frac{\overrightarrow{\mathcal{P}}}{\overleftarrow{\mathcal{P}}}\,.$ (11)
For Markovian systems [such as Eq. (2)], it is helpful to decompose the trajectory probabilities using the infinitesimal-time propagator, $\Pr(\mathbf{x}_{k+1},\lambda_{k+1}|\mathbf{x}_{k},\lambda_{k})$, as $\overrightarrow{\mathcal{P}}=p(\mathbf{x}_{0},0)\prod_{k=1}^{K}\Pr(\mathbf{x}_{k},\lambda_{k}|\mathbf{x}_{k-1},\lambda_{k-1})$. This decomposes the IEP rate at time $t=k\Delta t$ into
$\dot{\Sigma}_{t}$ $=$ $\dot{S}_{t}+\dot{\Omega}_{t}\,,$
$\dot{\Omega}_{t}$ $=$ $\lim_{\Delta t\to 0}\frac{1}{\Delta t}\ln\frac{\Pr(\mathbf{x}_{k+1}\,,\lambda_{k+1}|\mathbf{x}_{k},\lambda_{k})}{\Pr(\mathbf{x}_{k},\lambda_{k}|\mathbf{x}_{k+1},\lambda_{k+1})}\,.$ (12)
In Eq. (12) we have identified the time-derivative of the entropy, $\dot{S}_{t}=\ln[p(\mathbf{x}_{k+1},(k+1)\Delta t)/p(\mathbf{x}_{k},k\Delta t)]/\Delta t$, which is related to the Langevin system through Eq. (9). We shall refer to the remainder, $\Omega_{t}$, as the “informatic heat” (IH). The propagator in the denominator, $\Pr(\mathbf{x}_{k},\lambda_{k}|\mathbf{x}_{k+1},\lambda_{k+1})$, is the exact same propagator of Eq. (2), where the jump process is reverted [i.e., the process starts at $\mathbf{x}_{k+1}$ at time $t=k\Delta t$ and ends up at $\mathbf{x}_{k}$ at time $(k+1)\Delta t$ under the protocol value $\lambda_{k+1}$].
Focusing on the IEP has two advantages. First, its ensemble average $\langle\Sigma_{t}\rangle$ is a Kullback-Leibler divergence and thus guaranteed to be nonnegative [50]. Second, the dissipation exactly coincides with the IEP up to $k_{\mathrm{B}}T$, $\dot{R}_{t}=k_{\mathrm{B}}T\dot{\Sigma}_{t}$, in a remarkably wide range of scenarios [30]. In particular, one finds $\dot{Q}_{t}=-k_{\mathrm{B}}T\dot{\Omega}_{t}$ [45] (hence the term “informatic heat” for $\Omega_{t}$). For these cases, therefore, one does not have to resort to simplistic models for the dynamics and then employ Eq. (8) to obtain the thermodynamic heat rate. Instead, one can access the thermodynamic heat rate from the statistics of trajectories [Eqs. (11) and (12)] without any prior knowledge of the underlying equations of motion [52, 53]. Clearly, the correspondence between $\dot{\Omega}_{t},\dot{\Sigma}_{t}$ and $\dot{Q}_{t},R_{t}$ assumes the ER with its well-defined temperature.
More specifically, in Ref. [54] the IH rate is found to satisfy [55]
$\dot{\Omega}_{t}=[(\boldsymbol{\mu}\cdot\mathbf{F})\cdot\mathbf{D}^{-1}]\circ\dot{\mathbf{X}}_{t}.$ (13)
Relating this result to the thermodynamic heat rate, Eq. (7), for any force $\mathbf{F}$, evidently requires the ER. Therefore, if $\mathbf{D}$ and $\boldsymbol{\mu}$ are not related by a scalar, the informatic $\dot{\Sigma}_{t}$ is no longer related to the measurable $\dot{R}_{t}$, even on average. Consequently, the nonnegativity of $\langle\dot{R}_{t}\rangle$ (i.e., the second law) can no longer be established based on trajectory statistics. Below we propose a definition of nonequilibrium temperature that mends this discrepancy.
Nonequilibrium temperature.—We return to the second law, Eq. (10). Noting that it refers only to the first moments of Eqs. (8) and (9), we average over the noise realizations (while using $\langle\boldsymbol{\eta}_{t}\rangle=\mathbf{0}$), and over the instantaneous position, where $\langle[\cdot](\mathbf{X}_{t},\Lambda_{t})\rangle=\int{\text{d}}\mathbf{x}\,p(\mathbf{x},t)[\cdot](\mathbf{x},\Lambda_{t})$. Integration by parts then gives $\langle\dot{Q}_{t}\rangle=-\int{\text{d}}\mathbf{x}\,\mathbf{J}\cdot\mathbf{F}$ and $\langle\dot{S}_{t}\rangle=-\int{\text{d}}\mathbf{x}\,\mathbf{J}\cdot{\boldsymbol{\nabla}}\ln p$. Let us posit that the second law of thermodynamics, Eq. (10), still holds, but with a nonequilibrium temperature $\mathcal{T}$. Using the expressions for $\langle\dot{Q}_{t}\rangle$ and $\langle\dot{S}_{t}\rangle$, we express the average dissipation rate of Eq. (10) as
$\langle\dot{R}_{t}\rangle=k_{\mathrm{B}}\mathcal{T}\int{\text{d}}\mathbf{x}\,\frac{\mathbf{J}}{p}\cdot\mathbf{D}^{-1}\cdot\left(\frac{\mathbf{D}}{k_{\mathrm{B}}\mathcal{T}}\cdot\mathbf{F}p-\mathbf{D}\cdot{\boldsymbol{\nabla}}p\right).$ (14)
This separates out in parentheses a term reminiscent of the probability flux $\mathbf{J}$ of Eq. (4), which depends on our choice of $\mathcal{T}$.
We now impose the following physical conditions on Eq. (14): (i) To satisfy the second law, $\langle\dot{R}_{t}\rangle$ should be a nonnegative functional of $\mathbf{J}$. (ii) To agree with Onsager’s linear-response theory [56, 57], the right-hand side should reduce to a quadratic form of $\mathbf{J}$ near equilibrium and turn the second-law inequality into an equality for reversible processes ($\mathbf{J}\to\mathbf{0}$). (iii) The variable $\mathcal{T}$ should be a configuration-independent scalar. Recalling that the diffusivity is a positive-definite matrix, the simplest choice for the dissipation functional satisfying these three conditions is $\langle\dot{R}_{t}\rangle=k_{\mathrm{B}}\mathcal{T}\int{\text{d}}\mathbf{x}(\mathbf{J}\cdot\mathbf{D}^{-1}\cdot\mathbf{J})/p$ [58]. To complete our construction, we recover the second law of thermodynamics, $\langle\dot{R}_{t}\rangle=k_{\mathrm{B}}\mathcal{T}\langle\dot{S}_{t}\rangle-\langle\dot{Q}_{t}\rangle\geq 0$, if we chose the following drive-dependent nonequilibrium temperature:
$k_{\mathrm{B}}\mathcal{T}=\frac{\int{\text{d}}\mathbf{x}\,\mathbf{J}\cdot\mathbf{F}}{\int{\text{d}}\mathbf{x}\,\mathbf{J}\cdot\mathbf{D}^{-1}\cdot(\boldsymbol{\mu}\cdot\mathbf{F})}\,.$ (15)
Equation (15) is our central result. If the ER holds, such that $\mathbf{D}^{-1}\cdot\boldsymbol{\mu}=(k_{\mathrm{B}}T)^{-1}$, we readily get $\mathcal{T}=T$. Otherwise, the generalized temperature is an intricate quantity. Being drive-dependent, it is evidently less universal than the conventional temperature. This is to be expected for nonequilibrium systems with arbitrarily complex $\mathbf{J}$, $\mathbf{D}$, $\boldsymbol{\mu}$, and $\mathbf{F}$. Furthermore, $\mathcal{T}$ depends on time via the protocol $\Lambda_{t}$ affecting all these quantities. Even for a constant protocol $\Lambda_{t}=\mathrm{const}$, $\mathcal{T}$ will change in time due to the instantaneous distribution $p(\mathbf{x},t)$ entering $\mathbf{J}$. Indeed, nonequilibrium steady states generally have $\mathbf{J}\neq\mathbf{0}$, as nonconservative forces and drift create persistent fluxes across scales. Despite these caveats, Eq. (15) should play an experimentally meaningful role as a generalized temperature in nonequilibrium thermodynamics [19, 17, 29], as discussed below.
Consequences.—First, we reestablish the connection between the IH and heat and between the IEP and dissipation, i.e., between information and thermodynamics. Substituting Eq. (2) in Eq. (13), we find
$\dot{\Omega}_{t}$ $=$ $(\boldsymbol{\mu}\cdot\mathbf{F})\cdot\mathbf{D}^{-1}\cdot(\boldsymbol{\mu}\cdot\mathbf{F})+{\boldsymbol{\nabla}}\cdot(\boldsymbol{\mu}\cdot\mathbf{F})$ (16)
$+[(\boldsymbol{\mu}\cdot\mathbf{F})\cdot\mathbf{D}^{-1}\cdot\mathbf{b}]\cdot\boldsymbol{\eta}_{t}\,,$
which, combined with Eq. (9) and the relation $\dot{\Sigma}_{t}=\dot{S}_{t}+\dot{\Omega}_{t}$, gives the IEP rate as
$\dot{\Sigma}_{t}=\frac{\mathbf{J}}{p}\cdot\mathbf{D}^{-1}\cdot\frac{\mathbf{J}}{p}+\frac{2}{p}{\boldsymbol{\nabla}}\cdot\mathbf{J}+\left(\frac{\mathbf{J}}{p}\cdot\mathbf{D}^{-1}\cdot\mathbf{b}\right)\cdot\boldsymbol{\eta}_{t}\,.$ (17)
Upon taking the averages of Eqs. (16) and (17), we obtain
$k_{\mathrm{B}}\mathcal{T}=-\frac{\langle\dot{Q}_{t}\rangle}{\langle\dot{\Omega}_{t}\rangle}\,,$ (18)
and
$\langle\dot{R}_{t}\rangle=k_{\mathrm{B}}\mathcal{T}\langle\dot{\Sigma}_{t}\rangle=k_{\mathrm{B}}\mathcal{T}\int{\text{d}}\mathbf{x}\frac{\mathbf{J}\cdot\mathbf{D}^{-1}\cdot\mathbf{J}}{p}\,\geq\,0\,.$ (19)
Note that the correspondence holds for the means only and not for the full statistics, $\dot{Q}_{t}\neq-k_{\mathrm{B}}\mathcal{T}\dot{\Omega}_{t}$ and $\dot{R}_{t}\neq k_{\mathrm{B}}\mathcal{T}\dot{\Sigma}_{t}$. For example, the variances are different, $\langle({\text{d}}\Omega_{t})^{2}\rangle_{\mathrm{c}}=2{\text{d}}t\int(\boldsymbol{\mu}\cdot\mathbf{F})\cdot\mathbf{D}\cdot(\boldsymbol{\mu}\cdot\mathbf{F})p{\text{d}}\mathbf{x}$, whereas $\langle({\text{d}}Q_{t})^{2}\rangle_{\mathrm{c}}=2{\text{d}}t\int\mathbf{F}\cdot\mathbf{D}\cdot\mathbf{F}p{\text{d}}\mathbf{x}\neq(k_{\mathrm{B}}\mathcal{T})^{2}\langle({\text{d}}\Omega_{t})^{2}\rangle_{\mathrm{c}}$. (The subscript $\mathrm{c}$ denotes a cumulant.) Since the Jarzynski equality [35, 36] and the Crooks fluctuation theorem [37] require statistics beyond the mean, they are not restored by the generalized $\mathcal{T}$. Thus, without the ER, the various FRs [34] remain related only to informatic quantities (the IEP and IH) [41] and not to thermodynamics (dissipation and heat).
Another relation to be examined is the one between the heat of stochastic systems and the breaking of FDT. Such a connection was established by Harada and Sasa [59], implicitly assuming the ER and a configuration-independent scalar mobility. (We recall that the ER and FDT are equivalent only in equilibrium [60].) Connections found recently for various active systems between FDT breaking and the IH (e.g., Refs. [39, 40]) have similarly made these two strong assumptions. As discussed above, without these assumptions the IH and the thermodynamic heat are separate quantities. In addition, the FDT is violated [61]. It is then unclear which heat is connected to FDT breaking — the IH or the thermodynamic heat? We find that neither connection can be established in the general case [61], implying that the Harada-Sasa relation breaks down as well.
Nevertheless, generalized versions of the FDT and the Harada-Sasa relation can be derived in the absence of the ER. Here we outline these results; see the Supplementary Material (SM) for details [61]. Instead of considering the response to an impulsive linear potential gradient as was done in Ref. [59], we consider a perturbation of the force of the form $\delta\mathbf{F}(\mathbf{x},s)=-\delta h_{\mu}(s)\boldsymbol{\mu}^{-1}(\mathbf{x},s)\cdot\mathbf{D}(\mathbf{x},s)\cdot{\boldsymbol{\nabla}}x_{\mu}$ at time $s<t$, in response to the external perturbation $\delta\mathbf{h}(s)$. The difference between the response of the average force to such perturbation and its correlation with the position are found to be related to the thermodynamic heat [61],
$\langle\dot{Q}_{s}\rangle=\lim_{t\to s^{+}}\left(\frac{{\text{d}}}{{\text{d}}s}\langle\mathbf{F}(\mathbf{X}_{t},t)\cdot\mathbf{X}_{s}\rangle\right.\\ \left.-\frac{\delta}{\delta\mathbf{h}(s)}\cdot\langle\mathbf{F}(\mathbf{X}_{t},t)\rangle\right)\,.$ (20)
If, instead of $\mathbf{F}$, we consider the observable $\mathbf{D}^{-1}\cdot\boldsymbol{\mu}\cdot\mathbf{F}$, we obtain a similar relation for the IH,
$\langle\dot{\Omega}_{s}\rangle=\lim_{t\to s^{+}}\left(\frac{{\text{d}}}{{\text{d}}s}\langle[(\boldsymbol{\mu}\cdot\mathbf{F})\cdot\mathbf{D}](\mathbf{X}_{t},t)\cdot\mathbf{X}_{s}\rangle\right.\\ \left.-\frac{\delta}{\delta\mathbf{h}(s)}\cdot\langle[(\boldsymbol{\mu}\cdot\mathbf{F})\cdot\mathbf{D}](\mathbf{X}_{t},t)\rangle\right)\,.$ (21)
Thus the IH and thermodynamic heat are related to violations of a generalized FDT for different observables. If the ER holds, the two observables coincide and the classical FDT is recovered. With the additional assumption of a configuration-independent scalar mobility the Harada-Sasa relation is recovered as well [61].
We now revisit additional classical-thermodynamic results demonstrating the relevance of the generalized temperature $\mathcal{T}$. More details regarding these consequences are found in the SM [61]. (i) Consider a process that starts and ends at the same protocol value, $\Lambda_{t_{\mathrm{f}}}=\Lambda_{0}$, and then allowed to relax to steady state. Since $\oint_{0}^{\infty}{\text{d}}t\langle\dot{S}_{t}\rangle=0$, we obtain the Clausius inequality, $\oint_{0}^{\infty}{\text{d}}t\langle\dot{Q}_{t}\rangle/(k_{\mathrm{B}}\mathcal{T})=-\oint_{0}^{\infty}{\text{d}}t\langle\dot{\Sigma}_{t}\rangle\leq 0$ [61]. (ii) Construct a Carnot-like cyclic engine consisting of an “isothermal” process (with constant $\mathcal{T}_{\mathrm{H}}$ and $\langle Q_{\mathrm{in}}\rangle>0$), an “adiabatic” process (in which the system “cools” down to $\mathcal{T}_{\mathrm{C}}$), followed by another isothermal process (with constant $\mathcal{T}_{\mathrm{C}}$, and $\langle Q_{\mathrm{out}}\rangle<0$), and an adiabatic process (where the system reaches $\mathcal{T}_{\mathrm{H}}$). The efficiency of such an engine is found as $\langle-W_{\mathrm{cycle}}\rangle/\langle Q_{\mathrm{in}}\rangle=1-\mathcal{T}_{\mathrm{C}}/\mathcal{T}_{\mathrm{H}}+k_{\mathrm{B}}\mathcal{T}_{\mathrm{C}}\langle\Sigma_{\mathrm{cycle}}\rangle/\langle Q_{\mathrm{in}}\rangle$, where $\langle\Sigma_{\mathrm{cycle}}\rangle$ is the mean IEP during a cycle [61]. Thus, the mean IEP quantifies the cycle’s irreversibility compared to a generalized Carnot efficiency, $1-\mathcal{T}_{\mathrm{C}}/\mathcal{T}_{\mathrm{H}}$. (See also Ref. [62] in the context of nonequilibrium engines.) (iii) Define the nonequilibrium free energy, $A_{t}=H_{t}-k_{\mathrm{B}}\mathcal{T}S_{t}$. It is easy to show that it bounds the extractable work in “isothermal” transformations, $\langle-\dot{W}_{t}\rangle=-(\langle\dot{A}_{t}\rangle+k_{\mathrm{B}}\mathcal{T}\langle\dot{\Sigma}_{t}\rangle)\leq-\langle\dot{A}_{t}\rangle$ [61].
An interesting consequence of Eq. (15) relates to a particular class of nonequilibrium systems having multiple “temperatures”, e.g., a mixture of different types of particles, each of which is subjected to a different “thermostat” [63, 64]. These models may be relevant, for instance, for mixtures of passive and active particles [65, 66, 67, 68, 69]. They yield a temperature-like scalar which is a weighted average of the thermostats’ temperatures. Considering a diagonal but nonscalar $\mathbf{D}^{-1}\cdot\boldsymbol{\mu}$, with different inverse temperatures as its eigenvalues, we may view Eq. (15) as a generalization of such a weighted average. Perturbing different global parameters amounts to taking a perturbation $\mathbf{J}$ along different axes, thus picking up different eigenvalues (inverse temperatures) of $\mathbf{D}^{-1}\cdot\boldsymbol{\mu}$. We show a simplistic example in the SM [61].
Conclusion.—We have analyzed a generic nonequilibrium Langevin system [Eq. (2)] which violates the ER [Eq. (1)]. This describes ubiquitous scenarios in driven and active materials since the relation between dissipation (mobility) and fluctuations (diffusivity) is a particular consequence of the equilibrium Boltzmann distribution [15]. We have demonstrated the far-reaching consequences of ER violation, which may be summarized as the detachment of informatic quantities (IH, IEP) from their physical thermodynamic counterparts (heat, dissipation) and the resulting invalidation of fluctuation theorems which rely on these correspondences.
To restore the information-thermodynamics correspondence we have imposed the second law of thermodynamics directly and identified a generalized temperature $\mathcal{T}$ [Eq. (15)], which gives all the thermodynamically required properties: nonnegative dissipation, zero dissipation at equilibrium, linear thermodynamic stability, and the temperature being a macroscopic (integrated) property. The generalized temperature is model-independent — with a calorimetric measurement of the heat and path-sampling for the IH, one may find the system’s $\mathcal{T}$ with no required knowledge of the underlying dynamics. The variety of tools available for the estimation of IEP from trajectories (e.g., sampling marginal distributions [70, 71, 53] or the thermodynamic uncertainty relations [72, 73]) are particularly useful, since in our construction the average IEP remains connected to the average thermodynamic dissipation. The generalized temperature does not restore the various FRs [35, 37, 34, 59], which in the absence of the ER no longer dictate the statistics of thermodynamic properties [41].
Our choice of generalized temperature in Eq. (15) is not unique. We could have chosen $1/(k_{\mathrm{B}}\mathcal{T})=[\langle\dot{S}_{t}\rangle-f(\mathbf{J})]/\langle\dot{Q}_{t}\rangle$ with any function $f(\mathbf{J})$ satisfying $f(\mathbf{J})\sim J^{2}$ for small $J$ and $f(\mathbf{J}\neq\mathbf{0})>0$. All those temperatures will produce the second law and its consequences and coincide with the conventional temperature $T$ when the ER holds. Our definition of $\mathcal{T}$ stems from the simple quadratic form chosen for the dissipation $R_{t}$, Eq. (19). Beside its simplicity, this choice yields the natural and useful correspondence $\langle\dot{Q}_{t}\rangle=-k_{\mathrm{B}}\mathcal{T}\langle\dot{\Omega}_{t}\rangle$.
The temperature $\mathcal{T}$ as presented here is a property of a single closed system exchanging heat with an arbitrary bath. It dictates the flow of heat into or out of the system based on the change in the system’s entropy. Whether $\mathcal{T}$ functions as an effective temperature in other respects remains to be explored. For example, for two systems in contact (i.e., immersed in a bath that drives each of them out of equilibrium), heat should flow from the system with the higher $\mathcal{T}$ to the one with the lower $\mathcal{T}$. These additional attributes are left for a future study.
The applicability of $\mathcal{T}$ as a nonequilibrium temperature can be put to test in simple experimental scenarios. For example, one can make the mobility and diffusivity separately isotropic or anisotropic by controlling, respectively, the particle shape and athermal randomized forces. This will break the ER and allow the tuning of $\mathcal{T}$ according to Eq. (15). One can then study various manifestations of the second law based on $\mathcal{T}$ and its spatial variation [29].
Acknowledgements.
We greatly benefited from discussions with David Dean and Yael Roichman. T.M. acknowledges funding from the Israel Science Foundation (Grant No. 1356/22).